please help! im confused!!


(1/3)^(5x+2)=(1/9)^3
solving for x
i only got (1/3)^(5x+2)=1/729
i still cant figure out what x is

Answer by lwsshak3(6491)

(1/3)^(5x+2)=(1/9)^3
solving for x
3^-(5x+2)=1/9^3
3^-(5x+2=1/(3^2)^3=1/3^6=3^-6
-(5x+2=-6
5x+2=6
5x=4
x=4/5